You are looking at the car that is currently selling at $65,000. Classic Autos is offering free credit (i.e. no interest charged on the borrowed amount) on the car. You pay $10,000 down today and then pay the remaining balance at the end of 6years. Premium Motors next door does not offer credit, but will give you $20,000 off the list price if you pay cash now. Assume annual compounding with 9% discount rate. Which of the following statement is TRUE?

A.

Premium autos is offering a better deal since its PV of cost is approximately 500 lower than that of Premium Motors.

B.

Classic autos is offering a better deal since its PV of cost is approximately 2,200 lower than that of Premium Motors.

C.

Classic autos is offering a better deal since its PV of cost is approximately 600 lower than that of Premium Motors.

D.

Premium autos is offering a better deal since its PV of cost is approximately 2,400 lower than that of Premium Motors.
Answer #1
CLASSIC AUTOS
a Payment in year0 $10,000
b Payment in year6 $55,000 (65000-10000)
Discount rate =9%=0.09
c=b/(1.09^6) Present Value of payment in year6 $32,794.70
d=a+c Present Value of Total payment $42,794.70
PREMIUM MOTORS
a Payment in Year 0 $45,000 (65000-20000)
Present Value of Total payment $45,000
CLASSIC AUTOS IS A BETTER DEAL
Difference in PV =45000-42794.70 $2,205.30
ANSWER B
Classic autos is offering a better deal since its PV of cost is approximately 2,200 lower than that of Premium Motors.